OpenStudy (anonymous):

A right triangle has a hypotenuse of 153 centimeters and a leg of 72 centimeters. What is the length of the other leg?

OpenStudy (mathmate):

Use pythagoras theorem a^2+b^2=c^2 where c=hypotenuse=153 for this problem.